# Modeling Domain and Feedback Transitions for Cross-Domain Sequential Recommendation

### v1

@@ Introduction
$\left\{\begin{array}[]{l}\mathcal{L}_{\text{single }}^{A}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{A}(\mathbf{E}^{\prime}_{C}+\mathbf{E}^{\prime}_{A})\right)_{a_{|S_{A}|+1}},\\ \mathcal{L}_{\text{single }}^{B}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{B}(\mathbf{E}^{\prime}_{C}+\mathbf{E}^{\prime}_{B})\right)_{b_{|S_{B}|+1}},\\ \mathcal{L}_{\text{cross }}^{A}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{A}(\mathbf{E}^{\prime}_{C})\right)_{a_{|S_{A}|+1}},\\ \mathcal{L}_{\text{cross }}^{B}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{B}(\mathbf{E}^{\prime}_{C})\right)_{b_{|S_{B}|+1}}.\end{array}\right.$ (16)
Experiments
Dataset Douban
Domain Book Movie Music
Users $26,877$ $28,718$ $23,822$
Items $187,520$ $57,424$ $185,574$
Records $1,097,148$ $2,828,585$ $1,387,216$
Density $0.0218\%$ $0.1715\%$ $0.0314\%$
Avg. Rating $4.0391$ $3.8101$ $4.1749$
Table 1: Statistics of three domains on Douban.
@@
GRU4Rec $1.31$ $1.28$ $1.39$ $0.93$ $1.59$ $1.97$ $1.62$ $1.58$ $1.80$ $0.99$ $2.13$ $2.83$
SASRec $1.35$ $1.32$ $1.43$ $0.98$ $1.62$ $1.97$ $1.89$ $1.87$ $2.09$ $1.24$ $2.45$ $3.13$
SRGNN $1.40$ $1.38$ $1.48$ $1.00$ $1.70$ $2.00$ $1.69$ $1.63$ $1.84$ $1.11$ $2.12$ $2.78$
CoNet $1.35$ $1.35$ $1.48$ $0.96$ $1.72$ $2.11$ $1.62$ $1.55$ $1.79$ $1.01$ $2.07$ $2.83$
$\pi$-Net $1.42$ $1.44$ $1.61$ $0.91$ $1.93$ $2.48$ $1.62$ $1.58$ $1.85$ $0.90$ $2.22$ $3.08$
C2DSR $1.52$ $1.49$ $1.64$ $1.06$ $1.89$ $2.45$ $1.93$ $1.92$ $2.15$ $1.25$ $2.54$ $3.26$
TriCDR $\underline{1.55}$ $\underline{1.54}$ $\underline{1.68}$ $\underline{1.10}$ $\underline{1.95}$ $\underline{2.39}$ $\underline{2.00}$ $\underline{1.98}$ $\underline{2.23}$ $\underline{1.30}$ $\underline{2.62}$ $\underline{3.37}$
$\text{Transition}^{2}$ 1.63 1.62 1.77 1.15 2.04 2.51 2.32 2.33 2.59 1.53 3.05 3.80
Table 2: Experimental results (%) of the Book-Movie domains. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
@@
GRU4Rec $1.24$ $1.19$ $1.31$ $0.90$ $1.47$ $1.84$ $1.18$ $1.15$ $1.29$ $0.77$ $1.51$ $1.93$
SASRec $1.22$ $1.20$ $1.30$ $0.88$ $1.49$ $1.82$ $1.25$ $1.23$ $1.36$ $0.84$ $1.59$ $1.98$
SRGNN $1.24$ $1.22$ $1.34$ $0.91$ $1.52$ $1.86$ $1.28$ $1.25$ $1.39$ $0.86$ $1.61$ $2.05$
CoNet $1.30$ $1.31$ $1.44$ $0.90$ $1.69$ $2.11$ $1.23$ $1.21$ $1.39$ $0.77$ $1.64$ $2.18$
$\pi$-Net $1.28$ $1.26$ $\underline{1.45}$ $0.75$ $\underline{1.75}$ 2.35 $1.20$ $1.17$ $1.38$ $0.64$ $1.67$ $\underline{2.34}$
C2DSR $1.30$ $1.26$ $1.38$ $0.94$ $1.56$ $1.92$ $1.30$ $\underline{1.30}$ $\underline{1.43}$ $\underline{0.88}$ $\underline{1.69}$ $2.10$
TriCDR $\underline{1.34}$ $\underline{1.31}$ $1.43$ $\underline{0.95}$ $1.64$ $2.03$ $\underline{1.30}$ $1.28$ $1.42$ $0.86$ $1.67$ $2.10$
$\text{Transition}^{2}$ 1.40 1.39 1.53 0.96 1.80 $\underline{2.22}$ 1.41 1.40 1.57 0.90 1.87 2.38
@@
Table 3: Experimental results (%) of the Book-Music domains. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
Expermental Setting
Datasets
To verify the effectiveness of the proposed $\text{Transition}^{2}$, we conduct experiments on a public dataset from Douban [^2] https://www.douban.com/. We select two CDSR scenarios for experiments: “Book-Movie” and “Book-Music”. To ensure sequential constraints, we retain cross-domain interaction sequences that include at least 3 items from each domain within one year. Ratings greater than 3 are considered positive feedback, while ratings less than or equal to 3 are considered negative feedback. We preprocess the data into a sequential recommendation format and split it based on timestamp records: the first 80% serves as the training set, the next 10% as the validation set, and the final 10% as the test set. Statistics are provided in Table 1.
Baselines
In this section, We compare $\text{Transition}^{2}$ with three representative sequential recommendation (SR) baselines and five cross-domain SR baselines:
Sequential recommendation baselines:
  • GRU4Rec (Hidasi et al. 2015) uses GRUs to capture sequential patterns in session-based data for personalized recommendations.
  • SASRec (Kang and McAuley 2018) uses self-attention to capture long-range dependencies in user behavior for effective recommendations.
  • SRGNN (Wu et al. 2019) applies Graph Neural Networks (GNN) to session-based recommendation, representing user behaviors as graphs to capture complex relationships and improve accuracy.
Cross-domain sequential recommendation baselines:
  • CoNet (Hu, Zhang, and Yang 2018) models interactions in two domains using base networks and transitions information via a cross-network.
  • $\pi$-Net (Ma et al. 2019) introducing a novel gating recurrent module to model and transfer knowledge across different domains.
  • C2DSR (Cao et al. 2022) uses a graph neural network to leverage inter-domain co-occurrences and employs a contrastive infomax objective to transfer cross-domain preferences by maximizing mutual information.
  • TriCDR (Ma et al. 2024) uses triple cross-domain attention and contrastive learning to model comprehensive cross-domain correlations.
Evaluation Metrics
To ensure unbiased evaluation, we employ the leave-one-out method, consistent with methodologies used in previous studies (Ye, Li, and Yao 2023; Kang and McAuley 2018). Following Rendle’s approach (Krichene and Rendle 2020), each validation/test case is assessed with $1,000$ scores, comprising $999$ negative items and $1$ positive item. The Top-K recommendation performance across these $1,000$ ranking lists is evaluated using metrics such as MRR@$10$ (Mean Reciprocal Rank) (Voorhees et al. 1999), NDCG@$5$, $10$ (Normalized Discounted Cumulative Gain) (Järvelin and Kekäläinen 2002), and HR@$1$, $5$, $10$ (Hit Ratio) (Koren, Bell, and Volinsky 2009).
Implementation Details
Our algorithm is implemented in PyTorch (Paszke et al. 2019) . The embedding size ($D$) and mini-batch size ($N$) are fixed at $256$, with training epochs set to $100$ and dropout fixed at $0.2$. We employ Adam (Kingma and Ba 2014) as the optimizer for parameter updates. The $L_{2}$ regularizer coefficient is chosen from {$-41e\text{-}$ 4, $-55e\text{-}$ 5, $-51e\text{-}$ 5}, and the learning rate $lr$ is selected from {$-31e\text{-}$ 3, $-45e\text{-}$ 4, $-41e\text{-}$ 4}. In C2DSR, we vary the depth of the GNN $L$ from {$1$, $2$, $3$}, and adjust the harmonic factor $\lambda$ from $0.1$ to $0.9$ in increments of $0.1$. For SASRec-based models, we incorporate two single-head attention blocks and learnable position embeddings. The channel number is set to $5$ for $\pi$-net, consistent with the original paper (Ma et al. 2019). In our approach, $N_{H}$ for Head Number is chosen from {$4$, $8$, $12$, $16$}. Best-performing models are determined based on the highest Mean Reciprocal Rank (MRR) performance on the validation set, and their results are reported on the test set. The experimental setup includes Ubuntu 18.04 with an Intel Xeon Gold 5218 CPU (64 cores, 128 threads), 754GB RAM, and four NVIDIA GeForce RTX 3090 GPUs, each with 24GB of memory. The GPU driver version is 550.78, and the CUDA version is 12.4.
@@
w/o TAG $1.58$ $1.57$ $1.73$ $1.08$ $2.03$ $2.50$ $2.27$ $2.28$ $2.54$ $1.50$ $3.01$ $3.80$
w/o CTM $1.56$ $1.56$ $1.69$ $1.13$ $1.95$ $2.36$ $2.17$ $2.18$ $2.42$ $1.43$ $2.87$ $2.62$
only $\mathbf{M}_{1}$ $1.59$ $1.59$ $1.72$ $1.14$ $2.00$ $2.41$ $2.36$ $2.37$ $2.63$ $1.56$ $3.11$ $3.92$
only $\mathbf{M}_{2}$ $1.56$ $1.56$ $1.69$ $1.12$ $1.95$ $2.36$ $2.16$ $2.16$ $2.41$ $1.41$ $2.85$ $3.62$
only $\mathbf{M}_{3}$ $1.56$ $1.56$ $1.69$ $1.13$ $1.95$ $2.35$ $2.18$ $2.18$ $2.43$ $1.42$ $2.88$ $3.64$
only $\mathbf{M}_{4}$ $1.55$ $1.54$ $1.68$ $1.13$ $1.93$ $2.34$ $2.19$ $2.19$ $2.44$ $1.44$ $2.88$ $3.64$
only $\mathbf{M}_{f}$ $1.57$ $1.57$ $1.71$ $1.14$ $1.97$ $2.39$ $2.31$ $2.32$ $2.58$ $1.51$ $3.06$ $3.78$
only $\widetilde{\mathbf{M}}_{f}$ $1.54$ $1.54$ $1.65$ $1.13$ $1.90$ $2.26$ $2.29$ $2.31$ $2.54$ $1.54$ $3.01$ $3.75$
$\text{Transition}^{2}$ $1.63$ $1.62$ $1.77$ $1.15$ $2.04$ $2.51$ $2.32$ $2.33$ $2.59$ $1.53$ $3.05$ $3.80$
Table 4: Ablation studies (%) on the Book-Movie Domains.
Experimental Results and Analysis
Overall Performance

### v2

@@ 4.4.1. Domain Representation Alignment
(15) $\mathcal{L}_{\text{A-align}}=-\left[\log\frac{\exp\left(\mathbf{e}^{\prime\prime\top}_{A}\mathbf{e}^{\prime}_{A}/\tau\right)}{\sum_{\mathbf{e}^{\prime-}_{A}\in\mathcal{N}_{A}^{\prime}}\exp\left(\mathbf{e}^{\prime\prime\top}_{A}\mathbf{e}^{\prime-}_{A}/\tau\right)}+\log\frac{\exp\left(\mathbf{e}^{\prime\top}_{A}\mathbf{e}^{\prime\prime}_{A}/\tau\right)}{\sum_{\mathbf{e}^{\prime\prime-}_{A}\in\mathcal{N}_{A}^{\prime\prime}}\exp\left(\mathbf{e}^{\prime\top}_{A}\mathbf{e}^{\prime\prime-}_{A}/\tau\right)}\right],$
where $\tau$ is the temperature hyperparameter, $\mathcal{N}_{A}^{\prime}$ denotes other users’ histories within the same batch of $\mathbf{e}^{\prime}_{A}$ as negative samples, $\mathcal{N}_{A}^{\prime\prime}$ denotes other users’ histories within the same batch of $\mathbf{e}^{\prime\prime}_{A}$ as negative samples. Similarly, we obtain the loss $\mathcal{L}_{\text{B-align}}$ for domain B. The total alignment loss is as follows:
(16) $\mathcal{L}_{\text{align}}=\mathcal{L}_{\text{A-align}}+\mathcal{L}_{\text{B-align}}.$
4.4.2. Feedback Representation Contrast
For items with different feedback types, we enforce their representations to be pulled apart within the same sequence to enable the model to perceive their differences. Specifically, we treat pairs of positive-feedback items in each user sequence as positive samples, and negative-feedback items as negative samples. For a general representation $\mathbf{E}$ (applicable to $\mathbf{E}^{\prime}_{A}$, $\mathbf{E}^{\prime}_{B}$, and $\mathbf{E}^{\prime}_{C}$), we formulate the in-sequence triplet loss as follows:
(17) $\mathcal{L}_{\text{cont}}^{\mathbf{E}}=\sum_{i=1}^{|\mathcal{P}_{+}|}\max\Big{(}\|\mathbf{e}_{i}-\frac{1}{|\mathcal{P}_{+}|}\sum_{j=1}^{|\mathcal{P}_{+}|}\mathbf{e}_{j}\|_{2}^{2}$
$-\|\mathbf{e}_{i}-\frac{1}{|\mathcal{N}_{-}|}\sum_{k=1}^{|\mathcal{N}_{-}|}\mathbf{e}_{k}\|_{2}^{2}+\alpha,0\Big{)},$
where $\mathbf{e}_{i}$ represents the embedding vector of the $i$-th item in the sequence, $\mathcal{P}_{+}$ denotes the set of positive sample indexes and $\mathcal{N}_{-}$ denotes the set of negative sample indexes within the sequence. We compute this loss for $\mathbf{E}^{\prime}_{A}$, $\mathbf{E}^{\prime}_{B}$, and $\mathbf{E}^{\prime}_{C}$ respectively, and sum them up to obtain the contrastive loss $\mathcal{L}_{\text{cont}}$:
(18) $\mathcal{L}_{\text{cont}}=\mathcal{L}_{\text{cont}}^{\mathbf{E}^{\prime}_{A}}+\mathcal{L}_{\text{cont}}^{\mathbf{E}^{\prime}_{B}}+\mathcal{L}_{\text{cont}}^{\mathbf{E}^{\prime}_{C}}.$
By introducing the losses $\mathcal{L}_{\text{align}}$ and $\mathcal{L}_{\text{cont}}$, $\text{Transition}^{2}$  aligns these three sets of item representations and captures the differences among items with different feedback types within each user behavior sequence.
4.5. Model Training
For the final model training loss, we designed losses $\mathcal{L}_{\text{single }}^{A}$ and $\mathcal{L}_{\text{single }}^{B}$ for single-domain recommendation tasks and losses $\mathcal{L}_{\text{cross }}^{A}$ and $\mathcal{L}_{\text{cross }}^{B}$ for cross-domain recommendation tasks. These are then combined to obtain the total loss with $\mathcal{L}_{\text{align}}$ and $\mathcal{L}_{\text{cont}}$:
(19) $\mathcal{L}_{\text{total }}=\underbrace{\mathcal{L}_{\text{single }}^{A}+\mathcal{L}_{\text{single }}^{B}}_{\text{Single-Domain Loss }}+\underbrace{\mathcal{L}_{\text{cross }}^{A}+\mathcal{L}_{\text{cross }}^{B}}_{\text{Cross-Domain Loss }}+\mu_{1}\mathcal{L}_{\text{align}}+\mu_{2}\mathcal{L}_{\text{cont}},$
where $\mu_{1}$ and $\mu_{2}$ are hyper-parameters and the recommendation losses are all defined as the cross-entropy loss between the predictions and the ground truth:
(20) $\left\{\begin{array}[]{l}\mathcal{L}_{\text{single }}^{A}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{A}(\mathbf{E}^{\prime}_{C}+\mathbf{E}^{\prime}_{A})\right)_{a_{|S_{A}|+1}},\\ \mathcal{L}_{\text{single }}^{B}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{B}(\mathbf{E}^{\prime}_{C}+\mathbf{E}^{\prime}_{B})\right)_{b_{|S_{B}|+1}},\\ \mathcal{L}_{\text{cross }}^{A}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{A}(\mathbf{E}^{\prime}_{C})\right)_{a_{|S_{A}|+1}},\\ \mathcal{L}_{\text{cross }}^{B}=-\operatorname{log}\mathrm{Softmax}\left(\mathrm{MLP}_{B}(\mathbf{E}^{\prime}_{C})\right)_{b_{|S_{B}|+1}}.\end{array}\right.$
5. Experiments
Table 1. Statistics of three domains on Douban.
Domain #Users #Items #Records Density Avg. Rating
Book 26,877 187,520 1,097,148 0.0218% 4.0391
Movie 28,718 57,424 2,828,585 0.1715% 3.8101
Music 23,822 185,574 1,387,216 0.0314% 4.1749
Table 2. Statistics of two domains on Amazon.
Domain #Users #Items #Records Density Avg. Rating
Book 18,788 93,041 22,507,155 1.2876% 4.2958
Movie 12,807 29,243 4,607,046 1.2301% 4.1869
Table 3. Experimental results of the Book-Movie domains on Douban. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
@@ 5. Experiments
GRU4Rec $0.131$ $0.128$ $0.139$ $0.093$ $0.159$ $0.197$ $0.162$ $0.158$ $0.180$ $0.099$ $0.213$ $0.283$
SASRec $0.135$ $0.132$ $0.143$ $0.098$ $0.162$ $0.197$ $0.189$ $0.187$ $0.209$ $0.124$ $0.245$ $0.313$
SRGNN $0.140$ $0.138$ $0.148$ $0.100$ $0.170$ $0.200$ $0.169$ $0.163$ $0.184$ $0.111$ $0.212$ $0.278$
CoNet $0.135$ $0.135$ $0.148$ $0.096$ $0.172$ $0.211$ $0.162$ $0.155$ $0.179$ $0.101$ $0.207$ $0.283$
$\pi$-Net $0.142$ $0.144$ $0.161$ $0.091$ $0.193$ $0.248$ $0.162$ $0.158$ $0.185$ $0.090$ $0.222$ $0.308$
C2DSR $0.152$ $0.149$ $0.164$ $0.106$ $0.189$ $0.245$ $0.193$ $0.192$ $0.215$ $0.125$ $0.254$ $0.326$
TriCDR $0.155$ $0.154$ $0.168$ $0.110$ $0.195$ $0.239$ $0.200$ $0.198$ $0.223$ $0.130$ $0.262$ $0.337$
CA-CDSR $\underline{0.161}$ $\underline{0.163}$ $\underline{0.180}$ $\underline{0.115}$ $\underline{0.205}$ $\underline{0.259}$ $\underline{0.214}$ $\underline{0.210}$ $\underline{0.235}$ $\underline{0.143}$ $\underline{0.279}$ $\underline{0.359}$
$\text{Transition}^{2}$ 0.179 0.178 0.196 0.124 0.227 0.283 0.227 0.227 0.254 0.150 0.298 0.379
Table 4. Experimental results of the Book-Music domains on Douban. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
@@
GRU4Rec $0.124$ $0.119$ $0.131$ $0.090$ $0.147$ $0.184$ $0.118$ $0.115$ $0.129$ $0.077$ $0.151$ $0.193$
SASRec $0.122$ $0.120$ $0.130$ $0.088$ $0.149$ $0.182$ $0.125$ $0.123$ $0.136$ $0.084$ $0.159$ $0.198$
SRGNN $0.124$ $0.122$ $0.134$ $0.091$ $0.152$ $0.186$ $0.128$ $0.125$ $0.139$ $0.086$ $0.161$ $0.205$
CoNet $0.130$ $0.131$ $0.144$ $0.090$ $0.169$ $0.211$ $0.123$ $0.121$ $0.139$ $0.077$ $0.164$ $0.218$
$\pi$-Net $0.127$ $0.128$ $0.147$ $0.075$ $0.164$ $0.204$ $0.124$ $0.121$ $0.140$ $0.072$ $0.167$ $0.210$
C2DSR $0.130$ $0.126$ $0.138$ $0.094$ $0.156$ $0.192$ $0.130$ $0.130$ $0.143$ $0.088$ $0.169$ $0.210$
TriCDR $0.134$ $0.131$ $0.143$ $0.095$ $0.164$ $0.203$ $0.130$ $0.128$ $0.142$ $0.086$ $0.167$ $0.210$
CA-CDSR $\underline{0.140}$ $\underline{0.136}$ $\underline{0.151}$ $\underline{0.097}$ $\underline{0.177}$ $\underline{0.225}$ $\underline{0.134}$ $\underline{0.134}$ $\underline{0.147}$ $\underline{0.088}$ $\underline{0.175}$ $\underline{0.223}$
$\text{Transition}^{2}$ 0.147 0.144 0.160 0.099 0.186 0.235 0.141 0.138 0.155 0.090 0.183 0.235
@@
Table 5. Experimental results of the Book-Movie domains on Amazon. The best result is bolded and the runner-up is underlined. Improvements over the second-best methods are significant (t-test, p-value $<$ 0.05).
@@
GRU4Rec $0.085$ $0.082$ $0.085$ $0.078$ $0.080$ $0.095$ $0.101$ $0.099$ $0.105$ $0.065$ $0.131$ $0.163$
SASRec $0.081$ $0.079$ $0.082$ $0.073$ $0.085$ $0.092$ $0.107$ $0.103$ $0.116$ $0.071$ $0.135$ $0.172$
SRGNN $0.088$ $0.085$ $0.088$ $0.085$ $0.102$ $0.116$ $0.110$ $0.115$ $0.129$ $0.075$ $0.131$ $0.175$
CoNet $0.092$ $0.088$ $0.095$ $0.070$ $0.105$ $0.127$ $0.101$ $0.095$ $0.110$ $0.063$ $0.125$ $0.171$
$\pi$-Net $0.095$ $0.092$ $0.100$ $0.069$ $0.113$ $0.129$ $0.107$ $0.100$ $0.116$ $0.068$ $0.131$ $0.179$
C2DSR $0.089$ $0.086$ $0.092$ $0.075$ $0.097$ $0.114$ $0.112$ $0.107$ $0.121$ $0.077$ $0.136$ $0.177$
TriCDR $\underline{0.098}$ $\underline{0.094}$ $0.099$ $0.083$ $0.104$ $0.119$ $0.121$ $0.115$ $0.124$ $\underline{0.095}$ $0.134$ $0.163$
CA-CDSR $0.094$ $0.093$ $\underline{0.102}$ $\underline{0.085}$ $\underline{0.114}$ $\underline{0.129}$ $\underline{0.124}$ $\underline{0.123}$ $\underline{0.135}$ $0.087$ $\underline{0.143}$ $\underline{0.179}$
$\text{Transition}^{2}$ 0.106 0.104 0.120 0.094 0.113 0.139 0.136 0.134 0.145 0.102 0.161 0.194
Table 6. Ablation studies on the Book-Movie Domains on Douban.
Methods Book-domain recommendation Movie-domain recommendation
MRR $\mathrm{NDCG}$ $\mathrm{HR}$ MRR NDCG $\mathrm{HR}$
@$10$$5$$10$$1$$5$$10$$10$$5$$10$$1$$5$$10$
w/o TAG $0.163$ $0.162$ $0.175$ $0.110$ $0.207$ $0.260$ $0.213$ $0.214$ $0.230$ $0.140$ $0.281$ $0.259$
w/o CTM $0.161$ $0.161$ $0.171$ $0.108$ $0.197$ $0.249$ $0.209$ $0.208$ $0.222$ $0.135$ $0.274$ $0.252$
w/o $\mathcal{L}_{\text{align}}$ $0.165$ $0.164$ $0.176$ $0.111$ $0.205$ $0.256$ $0.210$ $0.211$ $0.218$ $0.141$ $0.285$ $0.265$
w/o $\mathcal{L}_{\text{cont}}$ $0.170$ $0.170$ $0.186$ $0.113$ $0.215$ $0.268$ $0.217$ $0.218$ $0.232$ $0.143$ $0.287$ $0.262$
only $\mathbf{M}_{1}$ $\underline{0.176}$ $\underline{0.176}$ $\underline{0.191}$ $\underline{0.122}$ $0.223$ $\underline{0.273}$ $\underline{0.226}$ $\underline{0.225}$ $\underline{0.252}$ $\underline{0.146}$ $\underline{0.291}$ $0.372$
only $\mathbf{M}_{2}$ $0.172$ $0.172$ $0.185$ $0.116$ $0.213$ $0.266$ $0.221$ $0.221$ $0.247$ $0.142$ $0.290$ $0.369$
only $\mathbf{M}_{3}$ $0.172$ $0.172$ $0.190$ $0.120$ $0.221$ $0.270$ $0.221$ $0.221$ $0.248$ $0.144$ $0.292$ $0.370$
only $\mathbf{M}_{4}$ $0.175$ $0.174$ $0.182$ $0.117$ $0.214$ $0.269$ $0.224$ $0.223$ $0.249$ $0.143$ $0.289$ $0.374$
only $\mathbf{M}_{f}$ $0.173$ $0.174$ $0.191$ $0.117$ $\underline{0.224}$ $0.270$ $0.224$ $0.223$ $0.241$ $0.139$ $0.281$ $\underline{0.375}$
only $\widetilde{\mathbf{M}}_{f}$ $0.172$ $0.172$ $0.189$ $0.116$ $0.216$ $0.261$ $0.222$ $0.222$ $0.235$ $0.142$ $0.284$ $0.373$
$\text{Transition}^{2}$ $\mathbf{0.179}$ $\mathbf{0.178}$ $\mathbf{0.196}$ $\mathbf{0.124}$ $\mathbf{0.227}$ $\mathbf{0.283}$ $\mathbf{0.227}$ $\mathbf{0.227}$ $\mathbf{0.254}$ $\mathbf{0.150}$ $\mathbf{0.298}$ $\mathbf{0.379}$
5.1. Expermental Setting
5.2. Datasets
To verify the effectiveness of the proposed $\text{Transition}^{2}$, we conduct experiments on two public dataset from Douban [^2] https://www.douban.com/ and Amazon [^3] https://jmcauley.ucsd.edu/data/amazon/. We select three CDSR scenarios for experiments: “Book-Movie” and “Book-Music” for Douban and “Book-Movie” for Amazon. To ensure sequential constraints, we retain cross-domain interaction sequences that include at least 3 items from each domain within one year. Ratings greater than 3 are considered positive feedback, while ratings less than or equal to 3 are considered negative feedback. We preprocess the data into a sequential recommendation format and split it based on timestamp records: the first 80% serves as the training set, the next 10% as the validation set, and the final 10% as the test set. Statistics are provided in Table 1 and Table 2.
5.3. Baselines
In this section, We compare $\text{Transition}^{2}$  with three representative sequential recommendation (SR) baselines and four cross-domain SR baselines:
Sequential recommendation baselines:
  • GRU4Rec (Hidasi et al., 2015) uses GRUs to capture sequential patterns in session-based data for personalized recommendations.
  • SASRec (Kang and McAuley, 2018) uses self-attention to capture long-range dependencies in user behavior for effective recommendations.
  • SRGNN (Wu et al., 2019) applies Graph Neural Networks (GNN) to session-based recommendation, representing user behaviors as graphs to capture complex relationships and improve accuracy.
Cross-domain sequential recommendation baselines:
  • CoNet (Hu et al., 2018) models interactions in two domains using base networks and transitions information via a cross-network.
  • $\pi$-Net (Ma et al., 2019) introducing a novel gating recurrent module to model and transfer knowledge across different domains.
  • C2DSR (Cao et al., 2022) uses a GNN to leverage inter-domain co-occurrences and employs a contrastive infomax objective to transfer cross-domain preferences by maximizing mutual information.
  • TriCDR (Ma et al., 2024) uses triple cross-domain attention and contrastive learning to model comprehensive cross-domain correlations.
  • CA-CDSR (Yin et al., 2024) focus on item representation alignment via sequence-aware augmentation and adaptive partial alignment.
5.4. Evaluation Metrics
To ensure unbiased evaluation, we employ the leave-one-out method, consistent with methodologies used in previous studies (Ye et al., 2023; Kang and McAuley, 2018; Zhang et al., 2025b; Shi et al., 2025). Following Rendle’s approach (Krichene and Rendle, 2020), each validation/test case is assessed with $1,000$ scores, comprising $999$ negative items and $1$ positive item. The Top-K recommendation performance across these $1,000$ ranking lists is evaluated using metrics such as MRR@$10$ (Mean Reciprocal Rank) (Voorhees et al., 1999), NDCG@$5$, $10$ (Normalized Discounted Cumulative Gain) (Järvelin and Kekäläinen, 2002), and HR@$1$, $5$, $10$ (Hit Ratio) (Koren et al., 2009), which are common evaluation metrics in recommendation (Zhang et al., 2024d; Chen et al., 2023; Wang et al., 2024; Shi et al., 2024; Zhang et al., 2024c).
5.5. Implementation Details
Our algorithm is implemented in PyTorch (Paszke et al., 2019) . The embedding size ($D$) and mini-batch size ($N$) are fixed at $256$, with training epochs set to $100$ and dropout fixed at $0.2$. We employ Adam (Kingma and Ba, 2014) as the optimizer for parameter updates. The $L_{2}$ regularizer coefficient is chosen from {$-41e\text{-}$ 4, $-55e\text{-}$ 5, $-51e\text{-}$ 5}, and the learning rate $lr$ is selected from {$-31e\text{-}$ 3, $-45e\text{-}$ 4, $-41e\text{-}$ 4}. In C2DSR, we vary the depth of the GNN $L$ from {$1$, $2$, $3$}, and adjust the harmonic factor $\lambda$ from $0.1$ to $0.9$ in increments of $0.1$. For SASRec-based models, we incorporate two single-head attention blocks and learnable position embeddings. The channel number is set to $5$ for $\pi$-net, consistent with the original paper (Ma et al., 2019). The magnitude of feature augmentation is fixed to $0.1$ in CA-CDSR. $N_{H}$ for Head Number is chosen from {$4$, $8$, $12$, $16$}. Best-performing models are determined based on the highest Mean Reciprocal Rank (MRR) performance on the validation set, and their results are reported on the test set. The experimental setup includes Ubuntu 18.04 with an Intel Xeon Gold 5218 CPU (64 cores, 128 threads), 754GB RAM, and four NVIDIA GeForce RTX 3090 GPUs, each with 24GB of memory. The GPU driver version is 550.78, and the CUDA version is 12.4.
5.6. Overall Performance